OpenStudy (anonymous):

If y varies directly as x, and y is 6 when x is 72, what is the value of y when x is 8?'

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

direct variation takes the form y = kx if y = 6 and x = 72, what is k?

OpenStudy (anonymous):

1/9!

Vocaloid (vocaloid):

not quite 6 = 72k what is k?

Vocaloid (vocaloid):

need to get some sleep, @dan815 would you mind taking over? I think they've got a few more questions after this one

OpenStudy (anonymous):

To find k, we need to get k by its self and we do that by dividing by 72 on each side of the equal sign. 6 = 72k 6 / 72 = 72k /72 6 / 72 = k 1 /12 = k k = 1 /12 Now that we found k (our constant of variation ) we can solve for "what is the value of y when x is 8" y = xk y = 8 * (1 / 12) y = 2/3